OpenStudy (anonymous):

solve the system using the substitution method .1x+.7y=-1.3 and .8x-.5y=7.9

OpenStudy (fibonaccichick666):

any ideas?

OpenStudy (anonymous):

need help

OpenStudy (fibonaccichick666):

I know, but I like to see if you have any thoughts first, so why don't we start by solving one of our eq for y

OpenStudy (fibonaccichick666):

can you do that?

OpenStudy (anonymous):

would start by doing this y=-(13/7)-x

OpenStudy (fibonaccichick666):

can you show mw all of your steps please?

OpenStudy (anonymous):

10(.1X+.7Y)=(-1.3)10 THEN would solve for y

OpenStudy (anonymous):

by subtracting x and divideing the 7

OpenStudy (fibonaccichick666):

when you divide by 7, do you divide the whole side or only part of it?

OpenStudy (anonymous):

the whole side so i would have y=-(13/7)-(1/7)x

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

try pluging it in i keep getting crazy answers

OpenStudy (fibonaccichick666):

so now, you are going to put that y=blahblah into your other eq and solve for x

OpenStudy (fibonaccichick666):

show me your steps again?

OpenStudy (anonymous):

8x-5(-(13/7)-(1/7))=79

OpenStudy (fibonaccichick666):

you forgot something

OpenStudy (anonymous):

ya the x

OpenStudy (fibonaccichick666):

yup so now, step by painfully slow step please :)

OpenStudy (anonymous):

8x+(65/7)+(5/7)x

OpenStudy (anonymous):

(61/7)x +(65/7)=79

OpenStudy (fibonaccichick666):

ok make sure you keep an = sign somewhere on your paper

OpenStudy (fibonaccichick666):

wait

OpenStudy (fibonaccichick666):

ok go (i'm really bad at 7*8 oops)

OpenStudy (anonymous):

(61/7)x=(488/70)

OpenStudy (anonymous):

x=8

OpenStudy (fibonaccichick666):

might I make a suggestion, get rid of the fractions first

OpenStudy (fibonaccichick666):

it'll make life easier

OpenStudy (anonymous):

yes so in the end i go x=8 and y=-3 is this correct

OpenStudy (fibonaccichick666):

looks like it let me check

OpenStudy (anonymous):

thanks for the help

OpenStudy (fibonaccichick666):

yup

OpenStudy (fibonaccichick666):

thats correct well done, just be careful with your distributions

OpenStudy (anonymous):

can you help me with this one as well m^2=19

OpenStudy (fibonaccichick666):

how do you undo a square?

OpenStudy (anonymous):

square root but this cannot be simplified into a whole number so the solution would be - square root of 19 and quire root of 19 correct

OpenStudy (fibonaccichick666):

mmhmm customarily we say \(\frac{+}{} \sqrt{19}\)

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

that's all there is to it

OpenStudy (anonymous):

thank you

OpenStudy (fibonaccichick666):

np :)